### v1

# Less is more: Probabilistic reduction is best explained by small-scale predictability measures

Abstract
The primary research questions of this paper center on defining the amount of context that is necessary and/or appropriate when investigating the relationship between language model probabilities and cognitive phenomena. We investigate whether whole utterances are necessary to observe probabilistic reduction and demonstrate that n-gram representations suffice as cognitive units of planning.
Less is more: Probabilistic reduction is best explained by small-scale predictability measures
Cassandra L. Jacobs Department of Linguistics University at Buffalo Buffalo, NY, USA cxjacobs@buffalo.edu          Andrés Buxó-Lugo Department of Psychology University at Buffalo Buffalo, NY, USA buxolugo@buffalo.edu
Anna K. Taylor EURECOM Université Paris - Sorbonne anna.taylor@eurecom.fr          Marie Leopold-Hooke Department of Computer Science and Engineering University at Buffalo Buffalo, NY, USA marieleo@buffalo.edu
1 Introduction
The articulatory dynamics of language production are sensitive to linguistic statistics, such as the frequencies of words, the phrases that those words appear in, and the relative probability of those words given the context dammalapati2021effects; bell2009predictability; harmon2021theory. Familiar words and phrases are assumed to be shorter due to factors like motor practice bybee2006usage, or other factors like the general ease of retrieval arnold2012audience. Previous studies have demonstrated that predictability influences articulatory dynamics, such that the more probable a word is, the shorter in duration the words tend to be lieberman1963some.
Words can be predictable in different ways. We can both anticipate the future, (e.g., "carve a…" is likely to end with pumpkin) as well as confirm that a word is supported by later information (e.g., pumpkin pie). Whereas the influence of word frequency on reduction could be explained by motor practice, findings that words are reduced when they are predictable has been argued to support rational production models. Conversely, lengthening improbable words is believed to afford listeners extra time to interpret the speech signal and minimize the odds of miscommunication aylett2004smooth; degen2020redundancy; futrell2023information; mahowald2013info; clark_relationship_2025; piantadosi2011word. However, probabilistic reduction effects are apparently illusory or contradictory depending on the level of granularity at which probabilities are computed (n-gram versus neural language model), and the type of word (content vs. function; bell2009predictability). Recent work has argued that large language models are uniquely suited to simultaneously explain probabilistic reduction and repetition reduction upadhye2025back; clark_relationship_2025. These studies have argued that n-grams are fundamentally unable to account for enough context to reliably model reduction relative to LLMs, which can encode up to 1024 tokens of context.
However, the cognitive plausibility of large language models as applied toward language production is unclear. Firstly, it is unknown what knowledge speakers draw on during language production, and how a speaker might efficiently estimate a word’s probability in context. The broad consensus in language production research is that speakers do not plan their utterances entirely in advance, and instead plan their utterances relatively incrementally ferreira2002incremental; futrell2023information; guhe2020incremental; pickering2013forward. So, the language production system might plan utterances at a lower level of granularity (e.g., one or two nearby words), or perhaps a larger one momma2021filling. Here we show that local probabilities suffice for the computation of probabilistic reduction, particularly when considering the computational efficiency of phrase-level retrieval arnon2013more; bannard2008stored.
To assess this possibility, the experiments described in this paper compare and contrast probability estimates of words derived from n-gram and LLM estimates, which are then incorporated as predictors in regression analyses of whole-word phonetic duration. We analyze four spontaneous speech corpora of American English: the Buckeye Corpus pitt2005buckeye, the Corpus of Regional African American Language (CORAAL; kendall2013speech; kendall2018corpus), Switchboard-NXT calhoun2010nxt of upadhye2025back, and the CANDOR corpus reece2023candor from clark_relationship_2025. Critically, our analyses include critical prosodic covariates that are known to affect word durations bishop2018anticipatory; baker2009variability (Section §2) and which we show to be partially confounded with language model probabilities (Section §5).
We demonstrate the utility of small-context modeling in Section §4 by showing that LLM probabilities extracted over intonational units, rather than whole conversational turns, better account for the effects of predictability on duration. We then show in Section §5 and §6 that statistics from large language models are not necessary to observe probabilistic reduction. In fact, the results of these analyses are broadly consistent with each other that LLM predictors produce worse model fits to durations than n-gram probabilities. We then examine the regression models across all datasets and observe that now-common practices for model comparison have masked the extremely small effect size of probabilistic factors on phonetic duration (as discussed in bell2009predictability).
2 Prosodic Preliminaries
A word’s duration (and, indeed, other dimensions of speech prosody) are sensitive to structural, probabilistic, and utterance-specific factors baker2009variability. In order to accurately gauge the contribution of probabilistic factors to duration, and determine whether large language models or multiword representations are more appropriate, we discuss these critical covariates.
2.1 Structural and lexical influences on duration
“Structural” factors that are known to affect duration include pauses, syntactic or prosodic phrase boundaries byrd2003elastic; klatt1975vowel; turk2007multiple; wightman1992segmental, in addition to lexical information such as number of phonemes and/or syllables, a word’s phonotactic probability goldrick2008phonotactic, and its phonological neighborhood density gahl2012reduce; gahl2016many.
2.2 Utterance length and position confounds on duration
Word durations have been shown to vary based on utterance planning/production dynamics. For example, words produced in isolation tend to have the longest durations as speakers may be disfluent and using timing information to hold the floor in a conversation clark2002using. Occasionally, researchers incorporate pause duration or pause location as a control variable clark_relationship_2025. This work instead conceptualizes phrase-final lengthening as a relative position effect, such that words closer to the end of a sentence tend to be longer.
Additionally, “anticipatory shortening” is a pattern where the durations of a syllable in a phrase are negatively correlated with the length of the phrase bishop2018anticipatory. This has been argued to reflect scope of planning effects, as syllable durations throughout the utterance are affected by a factor that is not made explicit until the production is complete. Thus, particularly for longer utterances, speech rates may be planned out somewhat in advance based on the approximate number of words or syllables a speaker plans. In contrast to prior work which has used speaker speech rate as a covariate, we use utterance-specific phrase length to account for gross timing patterns.
We determine utterances boundaries based on inter-pause units, which produce prosodic phrases, which have been argued to represent a fundamental unit of planning in production wheeldon2002minimal. Segmenting the utterance along inter-pause intervals and encoding a word’s relative position in an utterance produces the canonical anticipatory shortening effects across all datasets except Switchboard, which we show in Figure 1. All lengths of utterance show the standard phrase-final lengthening effect.
Figure 1: Sentence-final lengthening and anticipatory shortening effects (relative position) on duration across four spontaneous speech corpora. Only utterances comprised of ten words or fewer are visualized; all are analyzed.
2.3 Probabilistic factors
“Probabilistic” factors that affect reduction include a word’s frequency, its forward predictability, and its backward predictability clark_relationship_2025; bell2009predictability; upadhye2025back. Words that appear in frequent phrases are also typically reduced arnon2013more. In all cases, a word is expected to be reduced in duration when it is more probable in context. Additionally, a word’s informativity has a moderating effect, such that words that are typically unpredictable will tend to be less reduced seyfarth2014word.
Taking bigram-based conditional probabilities as an example, a word’s probability at position $i$ in context can be broadly generalized as $\frac{p(w_{i-1},w_{i})}{p(w_{i})}$ for forward transition probability, and $\frac{p(w_{i},w_{i+1})}{p(w_{i+1})}$ for backward transition probabilities. Large language models typically incorporate more context than n-gram probabilities, and can be computed for very large sequences of text, such as an entire conversation clark_relationship_2025, but the computation is similar.
Predictor of interest Transform Expected coefficient sign
Utterance length in number of words log -
Relative position of word within utterance inverse log +
Number of phonemes +
Forward predictability log -
Backward predictability log -
Table 1: Expected relationships of predictors on whole-word duration.
2.4 Covariate selection
To control for non-probabilistic factors that could influence phonetic duration, we extracted several predictor variables from previously established norming databases. For consistency with previous work, we use frequency information from the SUBTLEX-US corpus brysbaert2009moving. We obtain norms for a word’s length in number of phonemes from chee2020consistency. Number of syllables was not found to be a significant predictor in any analysis, and was recommended for exclusion in all model comparisons; therefore, we do not report any models with it included. Each word’s forward and backward transition probability is computed per dataset by the relevant n-gram or neural language model, described in each section. Table 1 shows the predicted outcomes from these analyses.
To account for phrase-final lengthening, we compute a transform on relative position, itself defined as the position ii111ii is strictly positive and non-zero. for a word $w_{i}$ out of all of the $n$ words in the sentence. We compute an inverse position value that produces higher values the closer the sentence is to the end.
$\text{Inverse relative position}=-\log(1-\frac{i}{n})$ (1)
At the end of the sentence, this value is undefined; therefore, we impute the maximum value from other non-initial positions in the other sentences in each corpus to determine this inverse position value.
We include random intercepts for each word in a particular part of speech (POS) category (e.g., extract as a noun or verb) to account for variation that might arise for homographic words or words used in different sentences as well as random intercepts by talker. For analysis of Buckeye and CORAAL in Section §6, course-grained POS tags were obtained through spaCy Honnibal_spaCy_Industrial-strength_Natural_2020. We relied on the Penn Treebank POS tags in Section §5, and no POS labels were available for Section §4.
3 Cognitive plausibility of backward language modeling
For speakers to know that a particular word will be easily understood because it is partly redundant with the downstream context, and can therefore be reduced aylett2004smooth; jaeger2010redundancy, the language production system must be able to compute such a probability efficiently.
The cognitive plausibility of LLMs relative to n-grams is less than obvious. It is widely believed that speakers store multiword sequences in long-term memory and that doing so promotes production fluency arnon2013more; bannard2008stored; morgan2024productive; siyanova2018production. By contrast, computing probabilities from a language model typically involves encoding the full, known upstream and/or downstream context to estimate a word’s probability. Alternately, some have applied the fill-in-the-middle procedure of bavarian2022efficienttraininglanguagemodels to more efficiently extract words’ probabilities, but this procedure does not clearly correspond to any known cognitive processes that speakers engage in to learn language, nor to perform fluent production. So, while the estimates from such a model may produce higher-quality predictors, the cognitive plausibility of these predictors is suspect. Finally, as upadhye2025back point out, solutions built on beam search are too expensive to be plausible accounts of production either meister-etal-2020-beam; pmlr-v97-cohen19a.
We propose that one efficient way for predictability to influence durations could be through the retrieval of multiword sequences (e.g., trigrams) from memory bannard2008stored; siyanova2018production. The subsequent experiments establish the feasibility of retrieving multiword sequences from memory as a way to explain probabilistic reduction.
4 Establishing short-range probabilistic reduction in clark_relationship_2025
In their analysis of the CANDOR corpus, clark_relationship_2025 found that words that are less predictable from the upstream context are produced with longer durations. As clark_relationship_2025 point out, the general length of a turn in conversation is rarely very long, typically under 50 words. Thus, their use of GPT-2’s full context window (up to 1024 tokens) to encode the entire turn is likely overly powerful for estimating a word’s probability. Thus, in this section, we re-analyzed the CANDOR corpus, but extracted language model probabilities only over intonational phrases using the probability estimation method of pimentel_how_2024.
As part of this reanalysis, we re-tokenized the utterances along pause boundaries to produce intonational phrases, which shortens the typical lengths of utterances substantially. We expect that smaller amounts of context will produce better estimates of a word’s probability in context, and therefore better predict the words’ durations.
We find that intonational phrase-level probabilities are more robust predictors of probabilistic production. Moreover, the two probability measures are highly correlated (Spearman $\rho$ = 0.69), and including both the short-term and long-term predictability measures hinders model convergence. Critically, the effect of predictability on duration is fully subsumed by the small-context estimate. The results of the regression model are included in Table 2. We visualize the stronger relationship to phonetic duration in Figure 2.
Transform Estimate Std. Error t value p
(Intercept) 0.254 0.034 7.455 <.001
SUBTLEX word frequency log -0.167 0.005 -28.980 <.001
Number of phonemes 0.142 0.003 55.654 <.001
Utterance length in words ($n$) log -0.055 0.003 -20.557 <.001
Short-range probability log -0.021 0.001 -38.126 <.001
Relative position inverse log 0.188 0.001 128.406 <.001
Table 2: Results of clark_relationship_2025 re-analysis.
Figure 2: Correlation between log language model probability and words’ durations (log ms) for short- and long-range language model probabilities.
5 N-gram versus language model predictability and probabilistic reduction in upadhye2025back
Probabilistic reduction may not apply as consistently across all words in the lexicon. Specifically, function words typically play a more structural or morphosyntactic role in English, and content words are more interchangeable. Consequently, some have claimed that function words are retrieved and planned at an earlier stage than content words garrett1975analysis. Consistent with this, bell2009predictability describe a dissociation in their analysis of Switchboard, in which content words were only reduced when backward transition probabilities were high, but function words were only reduced when forward transition probabilities were high.
An alternative position to the separate planning of content and function words is that the content-function distinction is somewhat fraught or arbitrary, or mostly reflects lexical frequency staub2024function. Similarly, if speakers retrieve phrase-shaped representations for production, then it is also plausible that probabilistic reduction should apply to both types of words equally. Such a finding would be consistent with upadhye2025back, whose results contradicted bell2009predictability, such that forward and backward predictability led to reduction on both content and function words in the Switchboard-NXT corpus. [^2] We wish to point out that the content-function distinction in the corpus is somewhat unclear. bell2009predictability state, ”The function-word category included discourse markers, conjunctions, existential there, pronouns and other proforms, prepositions, articles, quantifiers, demonstratives, verb auxiliaries, and verb particles. The remainder, nouns, verbs, adverbs, and adjectives, made up the content-word category.” However, an examination of the dataset reveals many words that belong to both categories, a large number of content words in the function word category, and pronouns notably included in the content word category. Future iterations of work in this vein should clarify or correct the content/function distinction. Therefore, in this experiment we replicate their analyses of reduction on content and function word durations.
The Switchboard-NXT data from upadhye2025back contains 650,570 observations. For language model probabilities, we use the provided autoregressive predictors from their custom GPT-2 model radford_language_2019 trained in both forward and backward directions on the CANDOR corpus333upadhye2025back present many other technological solutions that we find interesting, but these are not the focus of the present work.. For analyses of n-gram fluency, we rely on their smoothed bigram estimates derived from the CANDOR corpus, which serve a similar role to the bigram predictors used in bell2009predictability and seyfarth2014word. We broadly replicate all of their results with bigram statistics, such that we observe probabilistic reduction on both content and function words in both forward and backward directions (Table 3). Importantly, we also find that language model probabilities provide a worse fit to the duration data than bigram probabilities ($\Delta LL=2562$, $\chi^{2}=5123.6$, $p$ < .001).
One concern about these results (raised in Footnote 2) is that the presence of reduction for both Content and Function words could be partly due to errors in the coding of Content or Function word category labels, or due to the somewhat arbitrary distinction between the two. We thus conducted an additional analysis without the word class distinction and interaction terms and obtained roughly similar results, again finding that probabilistic reduction in both forward and backward directions is observable with simple bigram statistics. However, future work should determine the extent to which label noise negatively impacts statistical power to detect differences in patterns of probabilistic reduction across content and function words.
Transform Estimate Std. Error t value p
(Intercept) 0.138 0.022 6.096 <.001
SUBTLEX word frequency log -0.124 0.004 -32.227 <.001
Number of phonemes 0.133 0.002 86.337 <.001
Utterance length in words ($n$) log 0.008 0.001 7.047 <.001
Function words log -0.211 0.012 -17.076 < .001
Forward bigram probability (Content) log -0.007 0.001 -38.126 <.001
Backward bigram probability (Content) log -0.034 0.001 -52.036 <.001
Forward bigram probability (Function) log -0.004 0.001 -4.835 <.001
Backward bigram probability (Function) log -0.007 0.001 -7.833 <.001
Relative position inverse log 0.092 0.001 144.307 <.001
Table 3: Results of upadhye2025back re-analysis, excluding other.
One reason for the poor explanatory power of LLM probabilities might lie in an apparent confound between language model estimates and position within an utterance. We visualize several complex relationships between position, length, and duration in Figure 3. For language model probabilities, forward transition probabilities decrease substantially over positions, with the specific trajectory affected slightly by utterance length. Backward predictability furthermore shows a complex, highly non-linear pattern in which early, non-initial words and utterance-final words are assigned greater probabilities than utterance-medial words. Since as words approach the ends of the sentence, durations also increase, the predictive power of language model probabilities may be an artifact of a word’s position within an utterance. Whereas autoregressive language models may encode word position in a sentence more explicitly haviv-etal-2022-transformer, such confounds may be less present with the n-gram estimates as they do not explicitly encode position information, a possibility we explore in Section §6.
Thus, this analysis shows that not only do large language models lack both cognitive plausibility, relative to phrase-based representations that support successful language production arnon2013more; siyanova2018production, they also provide worse fits to the data, possibly due to major confounding factors.
Figure 3: Relationship utterance length and relative position on language model probabilities and duration in Switchboard corpus.
6 Examining probabilistic reduction in smaller and diverse dialect corpora
Previous work on probabilistic reduction has focused on White Mainstream American English (WMAE), such as that spoken in the Buckeye Corpus pitt2005buckeye and Switchboard godfrey1992switchboard; calhoun2010nxt. This work examines both WMAE and several varieties of African American Language (AAL), as AAL varieties are under-represented in computational psycholinguistics (as elsewhere in natural language processing; 10887767). Dialects represent a critical test case for modeling probabilistic language processing, because speech communities, and thus the linguistic statistics of those communities, differ along many dimensions.
A major challenge for assessing probabilistic reduction in dialect corpora, and smaller datasets more broadly, is that large language models may not produce high-quality estimates, either when trained on the corpora directly, or when fine-tuned. Moreover, while there is no reason to expect that predictability would not shape production in a similar way across dialects, if dialects are modeled by variables extracted from mainstream varieties (e.g., the models used in clark_relationship_2025 and upadhye2025back), then these biases are expected to distort any effects of linguistic statistical knowledge on production. This experiment therefore explores the feasibility of applying n-gram statistics to probabilistic reduction across smaller and diverse dialect corpora.
It is important to level the playing field so that LLM probabilities can capture dialect-specific variability in a low-data training regime, particularly since n-gram models trained over a specific corpus necessarily capture variability intrinsic to a dialect. vskrjanec2026language showed that fine-tuning language models (rather than training a small-scale neural language model from scratch) provides one means to capture the knowledge of a person’s specific linguistic experience to explain what they will find difficult to process (or produce). Thus, our experiments in this section compare fine-tuned language models against n-gram models of words’ forward and backward predictability.
6.1 Data preprocessing
In our experiments, we compare and contrast two datasets of spontaneous speech corpora representing Mainstream American English and African American Language. For MAE, we use the Buckeye Corpus used in seyfarth2014word, which contains 282,565 tokens. We compare probabilistic reduction in Buckeye with CORAAL kendall2018corpus. The full CORAAL corpus is much larger than Buckeye, with 1.37 million tokens, and as of this writing is composed of eight distinct dialect groups from across the United States. Since only some dialects had time-aligned transcripts, we applied a CORAAL-specific language model in the Montreal Forced Aligner (MFA) to obtain durations for all recordings mcauliffe2017montreal [^4] The MFA language model was trained on only four dialects: DCA, DCB, PRV, and ROC..
As in Section §4, we tokenized the spoken corpora into utterances along pause boundaries, as these roughly correspond to intonational units bishop2018anticipatory. Buckeye and CORAAL defined pauses somewhat differently, with CORAAL using phonetic criteria (60-70 ms of silence; kendallthesis09; kendall2013speech) and Buckeye relying on the intuition of the annotators pitt2005buckeye; raymond2003analysis. Furthermore, AAL has many grammatical and lexical features that are relatively unfamiliar to the base model, relative to WMAE. So, for the CORAAL data, we merged each utterance with the next to augment the input size and ensure a strong baseline for both datasets.
6.2 Language model fine-tuning
As language model probabilities extracted over short sequences appear to be better predictors of word durations (see Section §4), we chose to fine-tune the Pythia-160M language model biderman2023emergent over inter-pause units for both Buckeye and CORAAL experiments. Pythia is an autoregressive language model with a high degree of psychometric predictive power, and it has been successfully applied to other psychophysical data such as human reading times oh2023transformer; oh2024frequency. Fine-tuning proceeded for both Buckeye and CORAAL models for 40 epochs. We verified convergence by visually inspecting validation set loss using Tensorboard tensorflow2015-whitepaper. We share additional fine-tuning parameters in the Appendix.
6.3 Reverse language model fine-tuning
To obtain an estimate of a word’s backward transition probability, we again start with Pythia-160M as a baseline model, and we separately fine-tune this model on reversed MAE and AAL corpora. We used the same fine-tuning parameters as for the forward-trained models.
The correlation between neural language model probabilities and relative position persists in these data. We visualize the relative lack of correlation, characterized primarily by confounds with utterance edges, that is obtained by a simple trigram model over the corpora in Figure 4. Specifically, utterance-initial words tend to be relatively improbable in trigram models. Forward probabilities typically increase in their predictability after one or two words to a consistent baseline. Backward probabilities similarly are relatively probable for the early words in an utterance, but these increase to relatively high probabilities by the beginning.
Figure 4: Relationship between utterance length and relative position on trigram conditional probabilities in CORAAL and Buckeye.
There is considerable variability across LLM and n-gram predictors in terms of whether they predict probabilistic reduction or not. Specifically, in the case of both CORAAL and Buckeye, LLM predictors sometimes predicted lengthening when words are more probable. On the other hand, all n-gram based predictors show probabilistic reduction. We visualize the model coefficients from analyses of both corpora and both predictor types in Figure 5. We respectively report the CORAAL and Buckeye regression analyses in Table 4 and Table 5 below.
In addition to inconsistent coefficient signs, LLM predictors also proved worse at accounting for variance in duration than trigram estimates ($\Delta LL_{\text{coraal}}$=1515, $\Delta LL_{\text{buckeye}}$=351). We note that all coefficients are very small, particularly when compared to other predictors in the model, a finding we return to in the next section.
Figure 5: Regression model coefficient estimates for probabilistic reduction by corpus, language model type, and direction.
Transform Estimate Std. Error t value p
(Intercept) 0.562 0.019 29.24 <.001
SUBTLEX word frequency log -0.132 0.003 -38.29 <.001
Number of phonemes 0.120 0.001 85.36 <.001
Utterance length in words ($n$) log -0.114 0.001 -117.33 <.001
Forward trigram probability log -0.021 0.0002 -67.37 <.001
Backward trigram probability log -0.004 0.0003 -15.72 <.001
Relative position inverse log 0.127 0.001 226.74 <.001
Table 4: Results of CORAAL analysis.
Transform Estimate Std. Error t value p
(Intercept) 0.811 0.029 28.30 <.001
SUBTLEX word frequency log -0.190 0.005 -35.99 <.001
Number of phonemes 0.129 0.002 59.42 <.001
Utterance length in words ($n$) log -0.100 0.002 -48.99 <.001
Forward trigram probability log -0.020 0.001 -31.62 <.001
Backward trigram probability log -0.007 0.001 -11.49 <.001
Relative position inverse log 0.121 0.001 95.30 <.001
Table 5: Results of Buckeye analysis.
We attempted to determine whether the different dialect groups in CORAAL varied in the extent to which probabilistic reduction was evident. A model that included dialect did not converge, and reduction effects were stable across all dialects. We visualize the stability of this pattern in Figure 6.
Figure 6: All dialects show the same effect of predictability on duration.
7 Effect size of probabilistic reduction
As we report, many previous studies have reported contradictory results about the relationship between a word’s predictability and its duration, including our own comparison between LLM and n-gram based predictors. In this section, we illustrate how the coefficient estimates of probabilistic reduction are dwarfed by the estimates of other factors, even though all probabilistic predictors are highly significant and model comparison strongly supports inclusion of these predictors in the full model.
One plausible explanation for inconsistencies across studies of probabilistic reduction effects likely lies in the absolute size of the effect. When considering coefficient $t$ values as a proxy for standardized effect sizes, many predictors, particularly “garden variety” prosodic ones such as the relative position of a word in an utterance, can be an order of magnitude larger than these contextual probability estimates. Figure 7 directly compares the results of the Switchboard re-analysis from Section §5 to the results of the CORAAL and Buckeye analysis.
While all n-gram analyses are consistent with each other in the sense that probabilistic reduction is observed across all experiments, the effects themselves could be easy to miss, and studies may be underpowered, particularly depending on regression model parameterization. This potential explanation is corroborated by the finding that LLM probabilities fail to consistently produce negative coefficients.
Figure 7: Coefficient estimate magnitudes of control variables and probabilistic predictors of interest.
[…]
8 Discussion
In this work, we demonstrated that phrase-based representations provide a strong baseline to identify probabilistic reduction in duration. Despite claims by clark_relationship_2025 and upadhye2025back that the power of large language models enables them to uniquely capture probabilistic reduction in forward and backward directions and the effect of givenness or discourse status, we find that such claims are emphatically not supported by the data. That is, once prosodic timing variables specific to whole utterances are considered, particularly word position and utterance length, LLM-based predictors provide worse fits to duration data, potentially due to a number of confounds between position, length, and language model probability estimates. Throughout all of our experiments, it is clear that less context produces higher-quality estimates of speech production data, and more consistently.
Such findings provide strong evidence, counter to the zeitgeist, that small-scale, relatively incremental production, is critical for capturing phonological encoding effects. We argue that the consistent explanatory power of phrase statistics, and smaller-scale large language model probability estimates, suggests that if speakers are in fact using some estimate of a word’s probability in context during the phonological encoding process, they are doing so in a relatively local way. What we observe as effects of rational production that reduce words that are highly predictable may instead reflect other processes cohenpriva26. For example, speakers may retrieve phrases and prosodic knowledge about those phrases from memory somewhat incrementally during production, which could lead to reduction that manifests as a function of probabilities, but which instead reflects something like bigram or trigram frequency bannard2008stored; arnon2013more.
In addition to being poor explainers of phonetic duration, the application of LLMs to the study of probabilistic reduction in general poses several ethical challenges. First, many language models are typically taken "off-the-shelf", without regard to the nature of the web origins of their training data. In addition to domain differences between spontaneous speech and written language, LLMs that superimpose the statistics of the dominant variety on marginalized varieties will ultimately perpetuate hegemonic data practices birhane2021impossibility; birhane2022values. Fine-tuning or training from scratch may serve as one mitigation strategy to reduce the biases caused by these decisions, but clearly this is not necessary or sufficient to show probabilistic reduction.
On a more practical note, the effect sizes of probabilistic that we demonstrated in this paper are so small as to be almost imperceptible. While the significance level or model comparison differences might appear large, the empirical effect sizes of probabilistic reduction suggest speedups of mere single-digit milliseconds between very probable and very improbable words. For the speaker, shortening words’ durations may provide some benefit in terms of resource conservation, but this link is less than clear, and further experimentation is needed to test this alternative explanation.
Appendix A Appendix
Fine-tuning parameters used in experiments in Section §6:
max_epochs 40
batch_size 64
betas (0.9, 0.98)
epsilon 1e-8
learning_rate 1e-4
lr_decay_steps 1048567
warmup_steps 1024
weight_decay 1e-5
Table 6: Fine-tuning parameters used in all experiments